A sample of 25 concession stand purchases at the October 22 matinee of Bride of Chucky showed a mean purchase of $5.29 with a standard deviation of $3.02. For the October 26 evening showing of the same movie, for a sample of 25 purchases the mean was $5.12 with a standard deviation of $2.14. The means appear to be very close, but not the variances. At α = .05, is there a difference in variances? Show all steps clearly, including an illustration of the decision rule.
(Data are from a project by statistics students Kim Dyer, Amy Pease, and Lyndsey Smith.

Answer by stanbon(75887)

A sample of 25 concession stand purchases at the October 22 matinee of Bride of Chucky showed a mean purchase of $5.29 with a standard deviation of $3.02. For the October 26 evening showing of the same movie, for a sample of 25 purchases the mean was $5.12 with a standard deviation of $2.14. The means appear to be very close, but not the variances. At α = .05, is there a difference in variances? Show all steps clearly, including an illustration of the decision rule.
---------------------------
Using a 2-Sample FTest
Ho: s1^2 = s2^2
Ha: They are not equal
-----------------------
test statistic = 3.02^2/2.14^2 = 1.9915..
critical value 2.02677
Note: My F-chart provides this as a close approximation.
---------------------
p-value: 0.09811
------------------
Conclusion: Since p-value > 5%, fail to reject Ho.
-----------
The test does not support saying there is a difference between the variances.
----------------------------------
Note: An illustration of the decision rule is not possible on this site.
========================================================================
Use Google to find "testing two variances" to see a good example.
========================================================================
Cheers,
Stan H.